# Fractional Chern Insulator and Quantum Anomalous Hall Crystal in Twisted MoTe2

### v2

@@ IV Excitation Gap and Charge Continuum
To study the dynamical properties, we compute the band-resolved spectral functions $A_{mm}(\mathbb{k},\omega)\equiv-2{\rm Im}G_{mm}^{R}(\mathbb{k},\omega)$ of the CI state at $\nu=1$, the FCI state at $\nu=2/3$, and the QAHC state at $\nu=1/2$. $G_{mm}^{R}(\mathbb{k},\omega)$ is the retarded Green’s function and $m={\rm u},{\rm l}$ labels the upper/lower band (see Supplementary Sec. III [56] for details). The dynamical results reveal distinct features for these topological states, including the flattened band, charge excitation continuum, and band folding, providing valuable insights into their distinct nature.
@@
Lastly, for the QAHC phase at $\nu=1/2$, as shown in Fig. 5c, the lower band also splits into two sub-bands, with a charge gap $\Delta_{1/2}\approx 6.2$ meV. Unlike the charge-uniform FCI states, there exists shadow bands near, e.g., the $\bm{\Gamma}$ point in the lower sub-band. This characteristic feature suggests the occurrence of Brillouin zone folding, due to the presence of CO in the QAHC phase. Considering the CO wave vector $\bm{q}_{\rm CO}=(0,\pm\pi)$, as observed from $\rho(\bm{q})$ in Fig. 4f, we shift the spectra of the lower two sub-bands. As seen in Fig. 5(c), the spectral weights shift from the $Y$$M$ segment to $\Gamma$$Y$, resulting in two folded bands. The Fermi level lies within the gap between these two bands, with the lower folded band being significantly flatter than the bare band. This closely resembles the band structure of the $\nu=1$ CI state shown in Fig. 5a. Thus, our dynamical calculations further support that the QAHC state at $\nu=1/2$ represents a Chern insulator coexisting with charge order in an enlarged unit cell.
@@ V Characteristic Temperature Scales
Figure 6: Finite-temperature properties for various fillings. a The magnetic moment $m^{z}$ and b the local density of states $A_{\rm loc}(\omega=0)$ are displayed using contour plots. The black dashed line with diamonds shows the specific heat peaks at different fillings, which indicate the FM transition. Temperature evolution of c specific heat $c_{e}$, d magnet moment $m^{z}$, and e $A_{\rm loc}(\omega=0)$ at typical fillings. f Chemical potential $\mu$ and g inverse compressibility $\kappa^{-1}=\partial\mu/\partial\nu$ for various fillings $\nu$. The calculations are performed on a XC4×24×242424$\times$ 24$\times$ 24 × 24 × 2 and $\epsilon_{r}=15$.
@@
In Fig. 6, we present various finite-temperature properties calculated using the state-of-the-art finite-temperature tensor network method, tanTRG [66]. As shown in Fig. 6c, the specific heat curves $c_{e}=\frac{1}{N}\left(\frac{\partial E}{\partial T}\right)_{\nu}$ exhibit prominent peaks at certain critical temperatures $T_{c}$, for fillings $\nu=1$, $1/2$, $2/3$, indicating the occurrence of FM transitions. To identify the FM transitions, we compute the averaged magnetic moment $m^{z}=\frac{1}{2}\frac{\sum_{i}\expectationvalue{n_{i\uparrow}-n_{i\downarrow}}}{\sum_{i}\expectationvalue{n_{i\uparrow}+n_{i\downarrow}}}$, and show the results in Fig. 6d. As temperature decreases, the magnetic moment $m^{z}$ increases rapidly near $T_{c}$ and converges to a polarized value of 0.5 for $T<T_{c}$, evidencing a FM transition. In addition to the cases with $\nu=1,1/2,2/3$, we also present results for the $\nu=1/3$ case in Fig. 6. Here, $m^{z}$ displays an increasing trend (though it has yet to converge to saturation), suggesting that the FM transition occurs below the lowest temperature of 1.45 K attained in the simulations.
@@
For general fillings, we present a contour plot of $m^{z}$ in Fig. 6a. This plot reveals the FM phase transition with a Curie temperature $T_{c}\gtrsim 2$ K for $0.4\lesssim\nu\lesssim 1.05$, which closely aligns with the experimental observation that FM order can be observed at 1.6 K for $0.4\lesssim\nu\lesssim 1.2$ [8]. In Fig. 6a, the FM transition temperatures exhibit an overall decreasing trend as $\nu$ deviates from 1, except at $\nu=2/3$, where a dip is observed. This non-monotonic behavior is also qualitatively consistent with recent experimental findings [8]. Nevertheless, the dip position in Fig. 6a exhibits a quantitative discrepancy with experimental results, occurring at $\nu\simeq 2/3$ rather than the observed filling ($\nu\approx 0.75$). This suggests that either certain interactions present in the experimental system has been neglected in our real-space model, or additional parameter optimization — particularly the $\epsilon_{r}$, $\xi$, truncation length, or the system width $N_{y}$ — may be required for a better agreement.
Besides the FM transition temperature, there are two additional temperature/energy scales that we can compare with experiments. As presented in the $\mu$-$\nu$ relation in Fig. 6f, a charge gap $\Delta^{*}_{1}\approx 17.8$ meV can be estimated for $\nu=1$ based on the low-temperature calculations. This value is in good agreement with the experimental value of $14\pm 1$ meV [12], as well as to the single-particle gap $\Delta_{1}\approx 18.2$ meV estimated from our spectral function results in Fig. 5. The appearance of such a charge gap leads to incompressible states, as signaled by the peaks in the inverse compressibility, $\kappa^{-1}=\partial\mu/\partial\nu$. In Fig. 6g, we estimate $\kappa^{-1}$ from the $\mu$-$\nu$ relation and find the incompressible states at $\nu=1$, $2/3$, and $1/3$ indicated by the pronounced peaks, which correspond to the ground states of CI, FCI, and CO states. Remarkably, these incompressible states have indeed been experimentally observed also by recognizing the $\kappa^{-1}$ peaks [7].
The thermal activation of charge excitation constitutes a different temperature scale accessible in experiments. As the temperature increases from zero in an insulating state, there are thermally activated charge excitations, which give rise to increased longitudinal conductivity and the system behaves like a “metal”. We compute the local density of states at Fermi level, a quantity sensitively reflecting metallicity, via an imaginary-time approximation [73, 74, 66], i.e., $A_{\rm loc}(\omega=0)\approx\frac{2\beta}{N}\sum_{i,\sigma}\expectationvalue{c_{i\sigma}^{\dagger}(\beta/2)c_{i\sigma}}$, where $N$ is the number of unit cells and $\beta$ is the inverse temperature. To suppress boundary effects, we average over the bulk half of the system when computing $A_{\rm loc}(\omega=0)$.
In Fig. 6e, we find $A_{\rm loc}(\omega=0)$ initially increases and then drops with increasing temperature, exhibiting a maximum about $T^{*}\approx 30$ K for $\nu=1,2/3$ and 1/3. We refer to this as the (characteristic) thermal activation temperature. Notably, the observed $T^{*}\simeq 30~{}\mathrm{K}$ is approximately 7 times lower than the charge gap $\Delta_{1}\simeq 211.2~{}\mathrm{K}$ for $\nu=1$ and about 3 times lower than $\Delta_{2/3}\simeq 97.4~{}\mathrm{K}$ for $\nu=2/3$, highlighting a clear separation of energy scales. Our numerical results may explain the experimental observations, where the thermal activation gaps extracted from transport measurements are also roughly 5 and 3 times smaller than the corresponding charge gaps obtained from local magnetometry probe [12, 10]. Moreover, both energy scales, $\Delta_{1}$ and $T^{*}$, are in turn significantly greater than the FM transition temperature, $T_{c}$, implying that it is the FM transition temperature that restricts the observation of FCI and CI phases in experiments.
@@
In Fig. 6b, we show a contour plot of $A_{\rm loc}(\omega=0)$ as a function of filling and temperature. At low temperatures, $A_{\rm loc}(\omega=0)$ is significantly suppressed for $\nu=1$, 2/3, and 1/3, aligning with the insulating bulk states of the CI, FCI, and CO at these fillings, as demonstrated by DMRG calculations. A similar suppression is observed at $\nu=1/2$, albeit at a lower temperature, possibly since the parameter $\epsilon_{r}=15$ is near the boundary between the charge-gapped QAHC and CO phases (see Fig. 4a). On the other hand, in the small hole-doped regime to the left of $\nu=1$ line, $A_{\rm loc}(\omega=0)$ clearly shows finite values till low temperature, which indicates the presence of metallic behaviors.
VI Discussion
@@
In this study, we establish a unified theoretical framework to elucidate the intricate interplay between strong correlations and topology in tMoTe2. Directly incorporating multiple bands and band-mixing effects, our accurate tensor-network calculations based on a realistic model successfully explain the existence of FCI states and account for related experimental observations, and predict the presence of QAHC states. Furthermore, our studies address a wide range of experimental observations, bridging the gap in understanding both quantum dynamics and finite-temperature effects.
@@
Our high-precision results not only closely align with experimental data but also provide valuable insights to future studies. Through a detailed exploration, we identify the FMz order in the range $0.2\lesssim\nu\leq 1$, which covers the experimentally discovered CI state at $\nu=1$, and FCI states at fillings $\nu=2/3$, 3/5, 4/7, 4/9, and 5/9 in tMoTe2 [7, 8, 9, 10, 11, 12, 13]. Our ground-state calculations unveil a rich $\epsilon_{r}$-$\nu$ phase diagram featuring CI, FCI, QAHC, and CO states. For instance, at $\nu=1/3$ we observe a robust FCI phase for $15.5<\epsilon_{r}\leq 23$ and a GWC phase for $10<\epsilon_{r}\leq 15.5$, where the latter agrees with recent experimental results [7]. Notably, the $\epsilon_{r}$-driven quantum phase transitions between FCI and QAHC or CO phases — particularly their orders and universality classes — have yet to be clarified and warrants future investigation. The QAHC states with integer Hall conductivity $\sigma_{xy}=1$ at fractional fillings $\nu=1/2$, 2/3, 3/5, 4/5, and nearby incommensurate fillings also constitute a prediction remaining to be further explored.
@@
Our results underscore the crucial role of well-formed density structures in stabilizing the QAHC states. In experiments, the QAHC states have been identified in graphene-based systems with higher sample quality [39, 40, 41, 42, 43]. Recent experiments on tMoTe2 have revealed substantial real-space disorder through direct magnetic imaging [12], highlighting the need for high-quality samples to realize QAHCs in this system. Moreover, we conjecture that some of the experimentally observed FCI states at fillings such as $\nu=1/2$, 2/3, and 3/5 may originate from the “melting” of QAHC states, when the CO is disrupted. Thermal fluctuations may also drive a transition from QAHC to FCI states [40]. Thus the determination of full finite-temperature phase diagram of model [1] can clarify the topological Hall states in tMoTe2 and guide experiments.
@@
The influence of external fields on QAHCs also requires further study. QAHCs can be tuned into FCI states via applied current [40], perpendicular electric fields (displacement fields) [40, 41, 42, 43], and other experimental parameters. In particular, the displacement field, which introduces an onsite potential difference $u_{\rm D}$ between the two sublattices, is a commonly used tunable parameter in experiments and can drive a quantum phase transition from topological Hall states to trivial insulators in tMoTe2 [8, 7, 9, 10]. Thus, the effects of the displacement field on various states, including the CI at $\nu=1$, FCI at $\nu=2/3$ [75], and QAHC at $\nu=1/2$, remain to be explored. The insights gained from real-space models (Eqs. [1, 3, 4]) may offer valuable guidance also for studying other moiré systems with similar honeycomb superlattices [76, 77], such as twisted WSe2, AB-stacked MoTe2/WSe2, and AA-stacked bilayer graphene, etc. In addition, an external magnetic field can effectively suppress spin fluctuations, potentially influence both the ground-state and finite-temperature properties [78]. Furthermore, a broader range of experimentally tunable parameters [77] — such as twist angle, filling, dielectric constant, mechanical strain, optical excitation, and screening length — also remain to be investigated using the real-space approach introduced in the present study.
@@
Beyond the ground-state property, our many-body results reveal three characteristic energy scales for topological Hall states: the FM transition temperature ($T_{c}$), charge activation temperature ($T^{*}$), and the zero-temperature charge gap ($\Delta_{\nu}$). The observed charge gaps at low temperature have been found to be about 7 (3) times greater than the thermal activation temperatures for $\nu=1$ ($\nu=2/3$), which is consistent with experiments and resolves a key experimental contradiction [12, 10]. One possible explanation involves excitons as in-gap bound states. Their proliferation at finite temperature could induce strong fluctuations and renormalization effects, lowering the thermal activation threshold for excitations. Similar phenomena have been observed numerically in the topological Mott insulator [71, 79, 80] and bosonic FCI phases [81]. Further investigation into the collective excitations in the topological Hall states — such as excitons [77], magneto-rotons [45], possible graviton modes [82] and fractional excitations [83] — and the comparisons with spectroscopic measurements are left for future studies. Lastly, our method facilitates the exploration of exotic non-Abelian fractional excitations, particularly in the recently discovered fractional quantum spin Hall state [84].
@@
Note added.— In preparation of our manuscript, we note a new experimental work [85] that observed a compressible metallic region between $\nu=2/3$ to 1 that breaks spin and valley symmetry, and integer quantum anomalous Hall crystal near $\nu=2/3$ and $3/5$ fillings. These experimental observations are consistent with our numerical results here. The unconventional superconductivity observed in this system — proposed to be chiral superconductivity through weak-coupling approach on a continuum model [86] — is now ready to be investigated using precision many-body calculations on the real-space model established here.
Acknowledgements.

### v3

@@ IV Excitation gap and charge continuum
To study the dynamical properties, we compute the band-resolved spectral functions $A_{mm}(\mathbb{k},\omega)\equiv-2{\rm Im}G_{mm}^{R}(\mathbb{k},\omega)$ of the CI state at $\nu=1$, the FCI state at $\nu=2/3$, and the QAHC state at $\nu=1/2$. $G_{mm}^{R}(\mathbb{k},\omega)$ is the retarded Green’s function and $m={\rm u},{\rm l}$ labels the upper/lower band (see Supplementary section V online for details). The dynamical results reveal distinct features for these topological states, including the flattened band, charge excitation continuum, and band folding, providing valuable insights into their distinct nature.
@@
[…]
@@ V Characteristic temperature scales
In Fig. 6, we present various finite-temperature properties calculated using the state-of-the-art thermal tensor network method, tanTRG [62]. To characterize the moiré states, we compute the averaged magnetic moment $m^{z}=\frac{1}{2}\frac{\sum_{i}\expectationvalue{n_{i\uparrow}-n_{i\downarrow}}}{\sum_{i}\expectationvalue{n_{i\uparrow}+n_{i\downarrow}}}$ and the imaginary-time proxy $A_{\rm loc}(\omega=0)\approx\frac{2\beta}{N}\sum_{i,\sigma}\expectationvalue{c_{i\sigma}^{\dagger}(\beta/2)c_{i\sigma}}$, for the local density of states at the Fermi level  [73, 74, 62]. To reduce the boundary effects, we take averages over the bulk of the system. Contour plots of $m^{z}$ and $A_{\rm loc}(\omega=0)$ versus temperature $T$ and filling $\nu$ are presented in Fig. 6a and Fig. 6b, respectively.
@@
In Fig. 6c, we show the specific heat results $c_{e}=\frac{1}{N}\left(\frac{\partial E}{\partial T}\right)_{\nu}$ for certain fillings $\nu$. Pronounced peaks observed at characteristic temperatures $T_{c}$ for $\nu=1,1/2,$ and $2/3$ signal the occurrence of phase transitions. Correspondingly, as depicted in Fig. 6d, the magnetic moment $m^{z}$ increases sharply near $T_{c}$ upon lowering the temperature and approaches the saturated value 0.5 for $T<T_{c}$, indicating a FM transition. In addition to the cases with $\nu=1,1/2,2/3$, we also present results for the $\nu=1/3$ case. Here, $m^{z}$ displays an increasing trend (though it has yet to converge to saturation), suggesting that the FM transition occurs below the lowest temperature of 1.45 K attained in the simulations.
@@
As shown in Fig. 6a, the Curie temperature of the FM transition is $T_{c}\gtrsim 2$ K for $0.4\lesssim\nu\lesssim 1.05$, which closely aligns with the experimental observation that FM order can be observed at 1.6 K for $0.4\lesssim\nu\lesssim 1.2$ [8]. The FM transition temperatures exhibit an overall decreasing trend as $\nu$ deviates from 1, except at $\nu=2/3$, where a dip is observed. This non-monotonic behavior is also qualitatively consistent with recent experimental findings [8]. Nevertheless, a quantitative discrepancy exists between the theoretical dip position at $\nu\simeq 2/3$ in Fig. 6a and the experimental observation, where the minimum is found at a filling factor of $\nu\approx 0.75$. This suggests that either certain interactions present in the experimental system is missing in our real-space model, or the model parameters — particularly the $\epsilon_{\mathrm{r}}$, $\xi$, truncation length, or the system width $N_{y}$ — need to be further optimized for a better agreement.
@@
As temperature rises, even an insulating state can exhibit thermally activated charge excitations that enhance longitudinal conductivity, rendering a metal-like character to the system. This crossover in electrical transport behavior can be monitored through the local density of states at the Fermi level, $A_{\rm loc}(\omega=0)$. At low temperatures, $A_{\rm loc}(\omega=0)$ is significantly suppressed for $\nu=1$, 2/3, and 1/3, aligning with the insulating bulk states of the CI, FCI, and CO at these fillings, as have been demonstrated in DMRG calculations. A similar suppression is observed at $\nu=1/2$, albeit at a lower temperature, possibly since the parameter $\epsilon_{\mathrm{r}}=15$ is near the boundary between the QAHC and CO phases (see Fig. 4b). On the other hand, in the small hole-doped regime to the left of $\nu=1$ line, $A_{\rm loc}(\omega=0)$ clearly shows finite values till low temperature, which indicates the presence of metallic behaviors.
As shown in Fig. 6e, the maximum of $A_{\rm loc}(\omega=0)$ defines a characteristic thermal activation temperature $T^{*}\approx 30$ K for $\nu=1,2/3$ and $1/3$. Notably, the observed $T^{*}\simeq 30~\mathrm{K}$ is approximately 7 times lower than the charge gap $\Delta_{1}\simeq 211.2~\mathrm{K}$ for $\nu=1$ and about 3 times lower than $\Delta_{2/3}\simeq 97.4~\mathrm{K}$ for $\nu=2/3$, highlighting a clear separation of energy scales. Our results may explain the experimental observations, where the thermal activation gaps extracted from transport measurements are also roughly 5 and 3 times smaller than the corresponding charge gaps obtained from local magnetometry probe in experiments [12, 10].
Moreover, in Fig. 6f, we can estimate the charge gap from the calculated $\mu$-$\nu$ relation. A clear jump in the chemical potential appears for $\nu=1$ at low temperatures, yielding a gap of $\Delta^{*}_{1}\approx 17.8$ meV. This value is in good agreement with the experimental value of $14\pm 1$ meV [12], as well as to the single-particle gap $\Delta_{1}\approx 18.2$ meV estimated from our spectral function results in Fig. 5. The formation of a charge gap produces incompressible states, marked by peaks in the inverse compressibility, $\kappa^{-1}=\partial\mu/\partial\nu$. As shown in Fig. 6g, $\kappa^{-1}$ derived from the $\mu$-$\nu$ relation exhibits pronounced peaks at $\nu=1$, $2/3$, and $1/3$, corresponding to the CI, FCI, and CO ground states, respectively. Remarkably, these incompressible states have indeed been experimentally observed also by recognizing the $\kappa^{-1}$ peaks [7].
Overall, we conclude that both energy scales, $\Delta_{1}^{(*)}$ and $T^{*}$, are significantly greater than the FM transition temperature. This suggests that the experimental observation of FCI and CI phases is ultimately limited by the FM transition temperature.
VI Conclusion and Discussion
@@
In this study, we establish a moiré superlattice model and uniform framework to elucidate the intricate interplay between strong correlations and topology in tMoTe2. Directly incorporating multiple bands and band-mixing effects, our accurate tensor-network calculations based on a realistic model successfully explain the existence of FCI states and account for related experimental observations, and predict the presence of QAHC states. Furthermore, our studies address a wide range of experimental observations, and shed light on both quantum dynamics and finite-temperature effects in the moiré system.
@@
Our simulations align with recent experiments. Through a detailed exploration, we identify the FMz order in the range $0.2\lesssim\nu\leq 1$, which covers the CI state at $\nu=1$, and FCI states at fillings $\nu=2/3$, 3/5, 4/7, 4/9, and 5/9 discovered in tMoTe2 [7, 8, 9, 10, 11, 12, 13]. Our ground-state calculations unveil a rich $\epsilon_{\mathrm{r}}$-$\nu$ phase diagram, consisting of CI, FCI, QAHC, and CO phases. For instance, at $\nu=1/3$ we observe a robust FCI phase for $15.5<\epsilon_{\mathrm{r}}\leq 23$ and a GWC phase for $10<\epsilon_{\mathrm{r}}\leq 15.5$, and the presence of the latter agrees with recent experimental results [7]. Our simulations suggest that, to observe the FCI and FMz phases at $\nu=1/3$ in experiments, it would require lower temperatures ($T<1.45$ K) and a larger dielectric constant $\epsilon_{\mathrm{r}}$. Notably, nature of $\epsilon_{\mathrm{r}}$-driven quantum phase transitions between FCI and QAHC or CO phases have yet to be clarified and warrants future investigation. The QAHC states with integer Hall conductivity $\sigma_{xy}=1$ at fractional fillings $\nu=1/2$, 2/3, 3/5, 4/5, and nearby incommensurate fillings (e.g., $\nu\simeq 0.63$) also constitute a prediction remaining to be explored.
@@
Our results indicate that the disruption of well-formed density structures destabilizes the QAHC phase, leading to its suppression or transformation into other competing states. In experiments, the QAHC states have been identified in graphene-based systems with higher sample quality [40, 41, 42, 43, 44]. Recent experiments on tMoTe2 have revealed substantial real-space disorder through direct magnetic imaging [12], highlighting the need for high-quality samples to realize QAHCs in this system. Moreover, we conjecture that some of the experimentally observed FCI states at fillings such as $\nu=1/2$, 2/3, and 3/5 may originate from the “melting” of CO in the QAHC states — thermal fluctuations may drive a transition from QAHC to FCI states [41]. Therefore, constructing the complete finite-temperature phase diagram for model (Eq. (1)) would not only clarify the nature of the topological Hall states in tMoTe2 but also offer valuable guidance for future experimental work.
@@
The influence of external fields on QAHCs also calls for further studies. QAHCs can be tuned into FCI states via applied current [41], perpendicular electric fields (displacement fields) [41, 42, 43, 44], and other experimental parameters. In particular, the displacement field, which introduces an onsite potential difference $u_{\rm D}$ between the two sublattices, is a commonly used tunable parameter in experiments and can drive a quantum phase transition from topological Hall states to trivial insulators in tMoTe2 [8, 7, 9, 10]. Thus, the effects of the displacement field on various states, including the CI at $\nu=1$, FCI at $\nu=2/3$ [75], and QAHC at $\nu=1/2$, remain to be explored. The insights gained from real-space models Eq. (1) may offer valuable guidance also for studying other moiré systems with similar honeycomb superlattices [76, 77], such as twisted WSe2, AB-stacked MoTe2/WSe2, and AA-stacked bilayer graphene, etc. In addition, an external magnetic field can effectively suppress spin fluctuations, potentially influence both the ground-state and finite-temperature properties [78]. Furthermore, a broader range of experimentally tunable parameters [77] — such as twist angle, filling, dielectric constant, mechanical strain, optical excitation, and screening length — also remain to be investigated using the real-space approach introduced in the present study.
@@
Beyond the ground-state property, our many-body results reveal three characteristic energy scales for topological Hall states: the FM transition temperature ($T_{c}$), charge activation temperature ($T^{*}$), and the zero-temperature charge gap ($\Delta_{\nu}$). The observed charge gaps at low temperature have been found to be about 7 (3) times greater than the thermal activation temperatures for $\nu=1$ ($\nu=2/3$), which is consistent with experiments and resolves a key experimental contradiction [12, 10]. One possible explanation for this observation attributes it to the formation of in-gap excitonic bound states. Their proliferation at finite temperature could induce strong fluctuations and renormalization effects, lowering the thermal activation threshold for excitations. Similar phenomena have been observed numerically in the topological Mott insulator [71, 79, 80] and bosonic FCI phases [81]. Further investigation into the collective excitations in the topological Hall states — such as excitons [77], magneto-rotons [46], possible graviton modes [82], and fractional excitations [83] — and the comparisons with spectroscopic measurements are left for future studies. Finally, the methodology presented here can, in principle, be used to explore exotic non-Abelian fractional excitations — particularly in the recently discovered fractional quantum spin Hall state [84].
@@
Note added.— During preparing our manuscript, we note an experimental work [85] that observed a compressible metallic region between $\nu=2/3$ to 1 that breaks spin and valley symmetry, as well as integer quantum anomalous Hall crystals near $\nu=2/3$ and $3/5$ fillings. These experimental observations are consistent with our theoretical results here. The unconventional superconductivity observed in this system — proposed to be chiral superconductivity through weak-coupling approach on a continuum model [86] — is now ready to be investigated using precision many-body calculations on the moiré superlattice model established here.
Conflict of interest
The authors declare that they have no conflict of interest.
Acknowledgments